OpenStudy (anonymous):

Find equations of the lines that map to themselves under the following transformation ...

OpenStudy (anonymous):

\[\left[\begin{matrix}2 & 0 & 0 \\ 0 & 1 & 0 \\ 0 & 0 & 1\end{matrix}\right]\left(\begin{matrix}x \\ y \\ z\end{matrix}\right) = \left(\begin{matrix}X \\ Y \\ Z\end{matrix}\right)\]

OpenStudy (anonymous):

Not sure the approache here; is it best to start from \[(x-a)/b = (y-c)/d = (z-e)/f\] and pick coordinates that are transformed for e.g. (a,b,c) maps to (2a,b,2z)

OpenStudy (anonymous):

and go from there?

OpenStudy (anonymous):

2x=X y=Y z=Z

OpenStudy (anonymous):

apologies that bottom right parameter should be 2

OpenStudy (anonymous):

then, 2z=Z :)

OpenStudy (anonymous):

2x=X y=Y 2z=Z is the image of x,y,x but how do you figure out the equation of the line that that maps to itself

OpenStudy (anonymous):

so for example i can see that the y axis is 'invariant' i.e. doesnt change so x=0,z=0 would be one of the correct answers for eg. 0,5,0 => 0,5,0

OpenStudy (anonymous):

there is more than one answer, and i'd like to see the technique used